OpenStudy (anonymous):

how do I find the integral of sin^3(x)cos^3(x)???

OpenStudy (zarkon):

\[\sin^3(x)\cos^3(x)=\sin^3(x)\cos^2(x)\cos(x)=\sin^3(x)[1-\sin^2(x)]\cos(x)\]